OpenStudy (anonymous):

2(2x-4)=3(2x-4) Solve for the variable

OpenStudy (anonymous):

no you distributed right but the second step would end up being 4=2x. which would mean x=2. if you plug in the value for the variable to check 4 doesn't work (8=12) but plugging in 2 gives you 0=0 which works.

OpenStudy (anonymous):

i think its \[2(2x-4)=3(2x-4)\] then 4x-8=6x-12 transpose 6x and 8.. 4x-6x=-12+8 -2x=-4 divide both by -2 so.. x=2